deathsindragon:

calvin received bimonthly paychecks of 2007.25 last year. if 17.1% of his yearly income got with held for federal income tax, how much got with held for federal income tax from each of calvins paychecks last year? A. $ 343. 23 B. $34.32 C. $82.38 D. $823.78

Hoodmemes:

(im using a caculator) So first I think u should solve his income monthly= \[2 \times 2007.25=4,014.50 dollars\] I'm on the edge of thinking u should multipy that 4,0.14 times the months in a year which is 12 it would look like this \[4,014.5 \times 12=48,174 \] Then divide 17.1% by 100 that looks like \[ 17.1\div 100=0.711\] now its a decimal U should next dooo multiplication 0.171 times 48,174 that looks like \[0.171 \times 48,174 =8237.754 \] So u would mean homie got payed 8237.754 last year all together? (I got help from my brother lol,thats as far as we could get with answers and stuff srry) If im not correct,somebody should tell me why etc that would really help me understand this more!

Hoodmemes:

fr can somebody tell me if im incorrect.

deathsindragon:

i dont think its right, its not in the answer list

Hoodmemes:

Darn it,welp i tried srry i couldnt help ya

deathsindragon:

i got it, it was a,lol

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @deathsindragon i got it, it was a,lol \(\color{#0cbb34}{\text{End of Quote}}\) what step did i miss or didnt include when i was tryna solve it?

Hoodmemes:

@Hoodmemes

Hoodmemes:

Funny how I answered this but still managed to get it wrong,till this day I still don't see my mistake.

Hoodmemes:

I see what I did wrong !!!!

Hoodmemes:

lemme answer it the right way rq

Hoodmemes:

The old way I answered questions U wanna divide 17.1 by 100 \[17.1\div100=0.171\] Now multiply that by 2007.25 (the amount of money in the bimonthly checks Calvin gets) \[0.171\times2007.25=343.23975\] (343.23975) can also be known as (343.23) so that was what makes answer option A correct lol